OpenStudy (anonymous):

use rational exponents to write single radical expression: ^4√8 * ^2√7 I have four ans to choose from: I came up with ^4√392

OpenStudy (anonymous):

hi jazil, did you copy the correct prob

OpenStudy (anonymous):

let me double check sorry fourth root √8 * second root √7

OpenStudy (anonymous):

recheck it now Mark o.

OpenStudy (anonymous):

\[\sqrt[4]{8} * \sqrt[2]{7}\]

OpenStudy (anonymous):

sqrt4 sqrt8 sqrt2 sqrt7

OpenStudy (anonymous):

=2*2sqrt2 (sqrt2 sqrt7) =2*2*2sqrt7 8sqrt7

OpenStudy (anonymous):

Mark thanks but what is the ans. look ahead at the format from the equation symbols

OpenStudy (anonymous):

sqrt4 sqrt8 sqrt2 sqrt7 =8sqrt7

OpenStudy (anonymous):

thats what i can read from the prob

OpenStudy (anonymous):

\[\sqrt[4]{8} * \sqrt[2]{7}\] or fourth root √8 * second root √7

OpenStudy (anonymous):

we are mult: I have four choices to choose from which are: 1. \[\sqrt[4]{175616}\] 2. \[\sqrt[8]{56}\] 3. \[\sqrt[4]{39}\] 4. \[1/392^{4}\]

OpenStudy (anonymous):

just like that or fourth root √8 * second root √7

OpenStudy (anonymous):

mark o are you still there

OpenStudy (anonymous):

fourth root √8 * second root √7 is this it?

OpenStudy (anonymous):

fourth root √8 * second root √7 =((8)^1/2)^1/4* ((7)^1/2)^1/2 (8)^1/8 * (7)^1/4

OpenStudy (anonymous):

mark o look at the four choices to choose from for the answer that I posted right before you and tell me which one is it

OpenStudy (anonymous):

sory i cant read the answers real good from here, just use the regular typing dont use equations

OpenStudy (anonymous):

i dont have equation editor here, cant download them bec of my low memory ofmy hard drive

OpenStudy (anonymous):

1. \[\sqrt[4]{175616}\] is this 175616(sqrt4)? 2. \[\sqrt[8]{56}\] 56sqrt[8]] ? 3. \[\sqrt[4]{39}\] 39sqrt4? 4. \[1/392^{4}\] 1/(392^4)?

OpenStudy (anonymous):

thats how i read them from here